can someone please help show me how to sovle the following problem? i^37
Answer by Alan3354(69443)

i^37
= i
--------
i^2 = -1
i^4 = 1
i^8 = 1
Any multiple of 4 in the exponent --> 1
--------------
37 = 9*4 + 1
Divide the exponent by 4 and use the remainder
i^37 = i^1
= i